# Distilling circularly polarized photon emission from magnetized chiral plasmas

### v1

@@ III.3 Nonzero quark-number and chiral chemical potentials $\mu$ and $\mu_{5}$
The numerical results for the overall degree of circular polarization in photon emission are presented in Fig. 5. The magnitude of the effect is comparable to that discussed in Sec. III.1, and the qualitative dependence on the strength of the magnetic field, temperature, and transverse momentum is similar. However, there is a significant difference too. The angular dependence is not symmetric under the reflection $\phi\to-\phi$. The degree of circular polarization is larger in magnitude in the direction opposite to the magnetic field. Of course, this is expected since the dominant left-handed emission is asymmetric.
@@
IV Discussion and Summary
In this study, we considered the effect of nonzero quark-number (electric charge) and chiral chemical potentials on the polarization of photon emission from a strongly magnetized QGP. In particular, we showed that the composition of circularly polarized photon emission changes in qualitative ways when the charge densities are nonzero.
@@ IV Discussion and Summary
It should be mention that the circularly polarized photon emission rates are symmetric with respect to the reflection in the transverse (reaction) plane when the electrical charge density is nonzero ($\mu\neq 0$), assuming the chiral charge density vanishes ($\mu_{5}=0$). However, this changes at nonzero chiral charge density ($\mu_{5}\neq 0$). When $\mu_{5}>0$, we show that the emission rate of right-handed (left-handed) circularly polarized photons is higher in the upper (lower) hemisphere. (Here the two hemispheres are defined by the magnetic field direction.) The roles of the two circular polarizations interchange when the sign of $\mu_{5}$ changes.
As expected, the combined effect of nonzero electrical and chiral charge densities in the QPG is given by a superposition of their individual effects. Generally, a nonzero electrical charge leads to an overall dominance of the photon emission with one circular polarization over the other. The additional chiral charge density produces spatial asymmetries (with respect to reflection in the transverse plane) in photon emission for each circular polarization. Moreover, while the rate for one circular polarization is higher in the upper hemisphere, the rate for the other circular polarization is higher in the lower hemisphere.
@@ Appendix A Matsubara sums
Appendix B Dirac traces
In the derivation of the photon polarization function, one has to calculate four different types of Dirac traces. For completeness, here we present the corresponding results:
@@ Appendix B Dirac traces
$T_{4,f}^{\mu\nu}(s_{1},s_{2})$ $=$ $\frac{1}{\ell_{f}^{4}}\,\mbox{tr}\left\{\gamma^{\mu}(\mathbf{r}_{\perp}\cdot\bm{\gamma}_{\perp})L_{n-1}^{1}\mathcal{P}_{5}^{s_{1}}\gamma^{\nu}(\mathbf{r}_{\perp}\cdot\bm{\gamma}_{\perp})L_{n^{\prime}-1}^{1}\mathcal{P}_{5}^{s_{2}}\right\},$ (23)
where we introduced the shorthand notation $\bar{p}_{\parallel}\gamma_{\parallel}=\bar{p}_{0}\gamma^{0}-p^{3}\gamma^{3}$ and $\bar{p}_{0}=p_{0}+\mu$. Also, for brevity of presentation, we omitted the explicit dependence of all functions on their arguments, i.e., $L_{n}^{\alpha}\equiv L_{n}^{\alpha}\left(\zeta\right)$, where $\zeta=r_{\perp}^{2}/(2\ell_{f}^{2})$.
Appendix C Polarization projections of the traces
Here we separate the results of traces in Eqs. (B) – (23) into individual polarization contributions. By using the definition in Eq. (2), we introduce the following polarization projections:
@@ Appendix C Polarization projections of the traces
$\mathcal{T}_{0}^{f}(s,s)$ $=$ $\left[(\bar{p}_{0}-\mu_{5}-p_{z}s_{\perp}s)(\bar{p}_{0}-k_{0}-s\mu_{5})+(p_{z}-k_{z})\left(p_{z}-s_{\perp}s\bar{p}_{0}+s_{\perp}\mu_{5}\right)\right]L_{n}L_{n^{\prime}}$ (28)
@@
Appendix D Integration over $r_{\perp}$
We have the following results for the integrals:
@@ Appendix D Integration over $r_{\perp}$
$=$ $2\pi\ell_{f}^{2}\left[(\bar{p}_{0}-\mu_{5}-p_{z}s_{\perp}s)(\bar{p}_{0}-k_{0}-s\mu_{5})+(p_{z}-k_{z})\left(p_{z}-s_{\perp}s\bar{p}_{0}+s_{\perp}\mu_{5}\right)\right]\,\mathcal{I}_{0,f}^{n,n^{\prime}}$
@@
$\mathcal{J}^{f}_{\parallel}(s,s)$ $=$ $\int d^{2}\mathbf{r}_{\perp}e^{-i\mathbf{r}_{\perp}\cdot\mathbf{k}_{\perp}}e^{-\mathbf{r}_{\perp}^{2}/(2\ell_{f}^{2})}\mathcal{T}_{\parallel}^{f}(s,s)=8\pi\,\mathcal{I}_{2,f}^{n-1,n^{\prime}-1}(\xi)-\mathcal{J}^{f}_{0}(s,s).$ (35)
@@
Appendix E Polarization projections of the imaginary part of $\Pi^{\mu\nu}(\Omega;\mathbf{k})$
The polarization projections of the imaginary part of $\Pi^{\mu\nu}(i\Omega_{m};\mathbf{k})$ are given by the following formal expressions:
@@ Appendix E Polarization projections of the imaginary part of $\Pi^{\mu\nu}(\Omega;\mathbf{k})$
By using Matsubara sum in Appendix A, we derive

### v2

@@ IV.3 Nonzero quark-number and chiral chemical potentials $\mu$ and $\mu_{5}$
The numerical results for the overall degree of circular polarization in photon emission are presented in Fig. 7. The magnitude of the effect is comparable to that discussed in Sec. IV.1, and the qualitative dependence on the strength of the magnetic field, temperature, and transverse momentum is similar. However, there is a significant difference too. The angular dependence is not symmetric under the reflection $\phi\to-\phi$. The degree of circular polarization is larger in magnitude in the direction opposite to the magnetic field. Of course, this is expected since the dominant left-handed emission is asymmetric.
@@
V Discussion and Summary
In this study, we investigated the effect of nonzero quark-number (electric charge) and chiral chemical potentials on the polarization of photon emission from a hot, strongly magnetized QGP. Under the assumption of sufficiently strong background magnetic fields and high temperatures, photon emission is dominated by leading-order one-to-two and two-to-one processes, as shown in Fig. 1, while contributions from gluon-mediated two-to-two processes are neglected as subleading.
This work extends previous studies in Refs. [31, 39, 40] that employed the same approximation but focused exclusively on unpolarized photon emission from QGP with vanishing chiral chemical potential. In contrast, here we demonstrate that the composition of circularly polarized photon emission undergoes qualitative changes when nonzero chiral charge density is introduced. More interestingly, the emission of left-handed and right-handed polarized photons can be partially separated (or “distilled”) into the regions above and below the reaction plane.
@@ V Discussion and Summary
It should be mentioned that the circularly polarized photon emission rates are symmetric with respect to the reflection in the transverse (reaction) plane when the electrical charge density is nonzero ($\mu\neq 0$), assuming the chiral charge density vanishes ($\mu_{5}=0$). However, this changes at nonzero chiral charge density ($\mu_{5}\neq 0$). When $\mu_{5}>0$, we show that the emission rate of right-handed (left-handed) circularly polarized photons is higher in the upper (lower) hemisphere. (Here the two hemispheres are defined by the magnetic field direction.) The roles of the two circular polarizations interchange when the sign of $\mu_{5}$ changes.
As expected, the combined effect of nonzero electrical and chiral charge densities in the QGP is given by a superposition of their individual effects. Generally, a nonzero electrical charge leads to an overall dominance of the photon emission with one circular polarization over the other. The additional chiral charge density produces spatial asymmetries (with respect to reflection in the transverse plane) in photon emission for each circular polarization. Moreover, while the rate for one circular polarization is higher in the upper hemisphere, the rate for the other circular polarization is higher in the lower hemisphere.
@@ Appendix A Matsubara sums
Appendix B Calculation of Dirac traces
In the derivation of the photon polarization function in Sec. III, one has to calculate four different types of Dirac traces, i.e.,
@@ Appendix B Calculation of Dirac traces
$T_{4,f}^{\mu\nu}(s_{1},s_{2})$ $=$ $\frac{1}{\ell_{f}^{4}}\,\mbox{tr}\left\{\gamma^{\mu}(\mathbf{r}_{\perp}\cdot\bm{\gamma}_{\perp})L_{n-1}^{1}\mathcal{P}_{5}^{s_{1}}\gamma^{\nu}(\mathbf{r}_{\perp}\cdot\bm{\gamma}_{\perp})L_{n^{\prime}-1}^{1}\mathcal{P}_{5}^{s_{2}}\right\}.$ (24)
Here we used the shorthand notation $\bar{p}_{\parallel}\gamma_{\parallel}=\bar{p}_{0}\gamma^{0}-p^{3}\gamma^{3}$ and $\bar{p}_{0}=p_{0}+\mu$. Also, for brevity of presentation, we omitted the explicit dependence of all functions on their arguments, i.e., $L_{n}^{\alpha}\equiv L_{n}^{\alpha}\left(\zeta\right)$, where $\zeta=r_{\perp}^{2}/(2\ell_{f}^{2})$.
For the purposes of this study, we must separate the results of traces in Eqs. (B) – (24) into individual polarization contributions. By using the definition in Eq. (2), we introduce the following polarization projections:
@@
$\mathcal{T}_{0}^{f}(s,s)$ $=$ $\left[(\bar{p}_{0}-s\mu_{5}-p_{z}s_{\perp}s)(\bar{p}_{0}-k_{0}-s\mu_{5})+(p_{z}-k_{z})\left(p_{z}-s_{\perp}s\bar{p}_{0}+s_{\perp}\mu_{5}\right)\right]L_{n}L_{n^{\prime}}$ (29)
@@
Appendix C Integration over transverse spatial coordinates ($\mathbf{r}_{\perp}$)
After substituting the results for the Dirac traces into the definition of the polarization tensor in Eq. (10), we have to integrate the resulting expressions over the transverse spatial coordinates. There corresponding integrals for the four different polarization components are given by:
@@ Appendix C Integration over transverse spatial coordinates ($\mathbf{r}_{\perp}$)
$=$ $2\pi\ell_{f}^{2}\left[(\bar{p}_{0}-s\mu_{5}-p_{z}s_{\perp}s)(\bar{p}_{0}-k_{0}-s\mu_{5})+(p_{z}-k_{z})\left(p_{z}-s_{\perp}s\bar{p}_{0}+s_{\perp}\mu_{5}\right)\right]\,\mathcal{I}_{0,f}^{n,n^{\prime}}(\xi)$
@@
$\mathcal{J}^{f}_{\parallel}(s,s)$ $=$ $\int d^{2}\mathbf{r}_{\perp}e^{-i\mathbf{r}_{\perp}\cdot\mathbf{k}_{\perp}}e^{-\mathbf{r}_{\perp}^{2}/(2\ell_{f}^{2})}\mathcal{T}_{\parallel}^{f}(s,s)=8\pi\,\mathcal{I}_{2,f}^{n-1,n^{\prime}-1}(\xi)-\mathcal{J}^{f}_{0}(s,s),$ (36)
@@
Appendix D Imaginary part of $\Pi^{\mu\nu}(\Omega;\mathbf{k})$
By making use of the results in Appendices B and C, we derive the following expressions for the polarization projections of the imaginary part of $\Pi^{\mu\nu}(i\Omega_{m};\mathbf{k})$:
@@ Appendix D Imaginary part of $\Pi^{\mu\nu}(\Omega;\mathbf{k})$
By using a general result for the Matsubara sum in Appendix A, we derive